OpenStudy (anonymous):

prove csc2x=cscxsecx/2

OpenStudy (lgbasallote):

do you happen to have a formula for csc (2x)?

OpenStudy (anonymous):

@igbasallote no i dont have a formula

OpenStudy (lgbasallote):

Lg lol...ig is another user

OpenStudy (anonymous):

@lgbasallote my bad

OpenStudy (lgbasallote):

poor ig haha

OpenStudy (anonymous):

What is the reciprocal of csc(x)? Use that to help solve this problem

OpenStudy (lgbasallote):

^what he said

OpenStudy (anonymous):

@lwilson i dont even know what the reciprocal is

OpenStudy (lgbasallote):

csc = 1/sin

OpenStudy (anonymous):

you know that csc2x=1/(sin2x), then sin2x=2(sinx)(cosx), so csc2x=1/(2sinxcosx). 1/sinx=cscx, and 1/cosx=secx. to sum up, csc2x=(cscx)(secx)/2

OpenStudy (anonymous):

now find an identity for sin(2x)

OpenStudy (lgbasallote):

\[\large \begin{align} & \csc (2x) = \frac{1}{\sin (2x)}\\\\ & \frac{1}{\sin (2x)} = \frac{1}{2\sin x \cos x}\\\\ & \frac{1}{2\sin x \cos x} = (\frac{1}{2})(\frac{1}{\sin x})(\frac{1}{\cos x}) \end{align} \]

OpenStudy (lgbasallote):

do you get that @kkayla

OpenStudy (anonymous):

@lgbasallote yes thanks